OpenStudy (anonymous):

Question 1.1. Which ordered pair makes the equation true? 2x – 5y = –10 (Points : 1) (5, 4) (–10, 4) (2, 5) (20, 4)

OpenStudy (anonymous):

I have very limited time to get this done..

OpenStudy (texaschic101):

this is just a matter of subbing your points in and seeing if it is true 2x - 5y = -10 (5,4)...x = 5 and y = 4 now sub into original equation 2(5) - 5(4) = -10 10 - 20 = -10 -10 = -10 (true-- this is one solution) (-10,4)...x = -10 and y = 4 now sub 2(-10) - 5(4) = -10 -20 - 20 = -10 -40 = -10 (incorrect--so this is not a solution I think you can probably do the rest from here....be sure and pay attention to your signs